# Non-Abelian anyons in a periodically-driven Abelian model

### v2

@@ Abstract
Floquet-Majorana modes. We verify that this Floquet modulation indeed induces topological properties, mimicking a $p+ip$ superconductor, and yields Floquet-Majorana modes. We study these properties by numerically computing the Floquet Bogoliubov-de Gennes (BdG) Hamiltonian $\hat{H}_{F,\mathrm{BdG}}$, defined by $\hat{H}_{F}=(1/2)(\hat{\bm{f}}^{\dagger},\hat{\bm{f}})\hat{H}_{F,\mathrm{BdG}}(\hat{\bm{f}},\hat{\bm{f}}^{\dagger})^{T}$, as well as its quasienergy spectrum and ground state parity, from the driven model of Eq. (3[25]. The topological phase is expected for $|\mu|<4J$ and is characterized by a non-zero Chern number of the negative-energy fermion band [16, 27, 26]. Tuning the frequency $\omega$ to this topological region ($\mu\simeq-2J$), in the absence of $e$ vortices [$\mathcal{S}_{p}^{(r)}=1$], we find four nearly-degenerate high-frequency ‘ground’ states on a torus, one in each superselection sector corresponding to periodic or anti-periodic boundary conditions in $x$ and $y$ direction [16]. The topological degeneracy is approached exponentially with increasing system size, as shown in Fig. (2)(a), and only three ground states have even fermion parity and are thus physical, signalling three Ising-anyon types $e,\mathds{1}_{\rm is},\psi$. On a cylinder, the quasienergy spectrum exhibits two edge modes [Fig. 2(b)], indicating the development of nontrivial topology and a non-zero Chern number.
@@
Introducing a pair of separated $e$ particles on the torus, approximate Majorana modes at near-zero quasienergy appear [Fig. 2(c)], which disappear when $\omega$ is detuned out of the topological phase. The high-frequency description of Eq. (5) thus correctly captures the topological properties of the system, which are not spoiled by the higher-order contributions to the high-frequency expansion neglected in (5). Increasing the driving amplitudes $J=\Delta$, while adapting $\mu=-2J$ to stay in the topological region, the fermions enter an anomalous Floquet phase [36, 37, 30]. Although the time-averaged description of Eq. (5) is no longer valid as they depart from the high-frequency limit, the system still exhibits Majorana modes, at both quasienergy zero and ±$\pm\pi/T$, see Fig. 2(c). This effect, only possible in driven systems [38, 39, 36, 40, 41, 42, 43, 44, 45], has been proposed as a means to realize non-Abelian braiding and quantum computation in 1D Kitaev quantum wires [46, 47, 48, 49].
@@
Non-Abelian exchange phases. We analyze whether the exchange statistics of two $e$ particles (vortices) in the fermion topological phase is non-Abelian, as predicted in the presence of Majorana modes [8, 28]. Two vortices are exchanged explicitly in different fusion sectors following Levin-Wen’s protocol [50, 8, 51], sketched in Fig. 2(d), and Ref. [24]. The protocol computes the difference in the Berry phase accumulated by moving vortices along two paths $P$ and $P^{\prime}$, sharing the same set of positions but involving a different order of vortex moves. This guarantees that the resulting phase is only determined by the exchange statistics [50]. For each intermediate vortex configuration |$\ket{\vec{n}_{i}^{e}}$ along the paths, we numerically determine the corresponding even-parity ground state |$\ket{\Phi_{i}}$ of the Floquet-BdG fermion Hamiltonian [25]. The Berry phase accumulated along the path $P$ is computed from the sequence of ground states as $\theta_{P}=\mathrm{Arg}\prod_{(i,i+1)\in P}\bra{\Phi_{i+1}}\hat{\tilde{S}}^{(r)}_{i+1,i}\ket{\Phi_{i}}$. Here, $\hat{\tilde{S}}^{(r)}_{i+1,i}$ is the fermionic part of the quasiparticle representation of the spin operator $\hat{Z}_{i}$ which converts the vortex configuration |$\ket{\vec{n}^{e}_{i}}$ into |$\ket{\vec{n}^{e}_{i+1}}$ by displacing one $e$ particle. Indeed, the Pauli operators $\hat{Z}_{v,R}$ and $\hat{Z}_{v,B}$ between vertices $v$ and $v+r$ (with $r=x$ or $y$), map to operators $\hat{t}_{v}^{(r)}=\hat{\tilde{\mathcal{S}}}_{v}^{(r)}(\hat{b}_{v}^{\dagger}+\hat{b}_{v})(\hat{b}_{v+r}^{\dagger}+\hat{b}_{v+r})$, describing boson motion but also containing a fermionic part to capture semion statistics: $\hat{\tilde{\mathcal{S}}}_{v}^{(x)}=1$ and $\hat{\tilde{\mathcal{S}}}_{v}^{(y)}=\prod_{p\in\mathrm{L}[p(v+y)]}(-1)^{\hat{f}_{p}^{\dagger}\hat{f}_{p}}$, where $\mathrm{L}[p(v+y)]$ denotes all plaquettes to the left of the spin between $v$ and $v+y$. Details about the numerical evaluation of the matrix elements in $\theta_{P}$ are given in the SM [25]. We compare the exchange phase in different fusion sectors by creating the $e$-particle pair on top of two different degenerate ground states. The prediction for Ising anyons is that the exchange phase $R_{\mathds{1}_{\rm is}}^{\sigma\sigma}$ in the $\mathds{1}_{\rm is}$ sector (where the vortices fuse to the vacuum) and $R_{\psi}^{\sigma\sigma}$ in the $\psi$ sector (where they fuse to a fermion $\psi$) differ by $\pi C/2$ for Chern number $C$, namely $R_{\mathds{1}_{\rm is}}^{\sigma\sigma}=e^{-i\pi{C}/8}$ and $R_{\psi}^{\sigma\sigma}=e^{3i\pi C/8}$ [8]— a signature of their non-Abelian nature [52]. The two sectors $\mathds{1}_{\rm is}$ and $\psi$ are selected by considering doubly-antiperiodic and doubly-periodic fermion boundary conditions, respectively [24]. In the antiperiodic case, the even-parity ground state is the quasiparticle vacuum, such that the pair of vortices is created from and fuses to the vacuum. For the periodic case, the BdG ground state has odd parity and thus the physical even-parity ground state is the first-excited state containing one BdG fermion. The vortex pair thus fuses to this fermion. The counterclockwise exchange phases obtained are shown in Fig 2(e), for varying driving amplitude in the high-frequency regime and for different system size. The results converge rapidly with the size towards the prediction for Ising topological order with $C=1$. Inverting the handedness of the circular drive sign, $\phi_{y}-\phi_{x}=-\pi/2$, inverts the sign of the exchange phases. In conclusion, we have shown that the $e$ vortices carrying Floquet-Majorana modes, arising in the driven toric code in the high-frequency regime, behave like non-Abelian Ising anyons.
Conclusion. With the example of the toric-code model, we have shown that time-periodic driving of an Abelian-anyon system can lead to the emergence of non-Abelian topological order. The ideas presented here may be extended to other 2D topologically-ordered systems, exploiting the fact that pointlike anyon quasiparticles can be driven through local control. Moreover, they indicate a potential path towards non-Abelian physics, building on out-of-equilibrium Abelian phases as recently realized in engineered quantum systems [6, 7, 5].
Acknowledgements.
@@
The author is grateful to André Eckardt, Nathan Harshman, Alexander Schnell, Isaac Tesfaye and Sandro Wimberger for helpful and inspiring discussions. The author also acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Research Unit FOR 2414 - project number No. 277974659.
@@
Non-Abelian anyons in a periodically-driven Abelian model
@@ I Quasiparticle mapping
where $\Gamma_{1,v}$ is a path connecting vertices 1 and $v$, while $\tilde{\Gamma}_{1,p}$ connects the $1$st and the $p$th plaquette, with the specific path convention depicted in Fig. S1(a): $\Gamma_{1,v}$ descends vertically from the top-left vertex until the vertex row to which $v$ belongs and then continues horizontally until $v$; $\tilde{\Gamma}_{1,p}$ connects the top-left plaquette to the right-most one in the same row, descends vertically until the plaquette row to which $p$ belongs, and then horizontally to reach $p$. With this choice, it can be verified that spin anti-commutation rules are satisfied by the pseudospins, $\{\hat{\bm{X}}_{1}^{e}\hat{\bm{X}}_{v}^{e},\hat{\bm{Z}}_{v}^{e}\}=0$ and $\{\hat{\bm{X}}_{1}^{\psi}\hat{\bm{X}}_{p}^{\psi},\hat{\bm{Z}}_{p}^{\psi}\}=0$. The operators in Eq. (I) are depicted in Fig. S2(a).
Using Eqs. (S6) and (I), the two-spin driving terms of the form $\hat{X}_{i}\hat{Z}_{j}$ entering the drive Hamiltonian $\hat{H}_{d}(t)$ of Eq. (2) are mapped to pseudospin operators. The precise mapping depends on whether a term involves spins located on the boundaries, which are intersected by $\Gamma_{r}$ or $\tilde{\Gamma}_{r}$, or not [24]. We only report, as an example, the mapping for terms in the bulk, not intersecting the boundary, which reads
@@
×$\times\Big{[}\prod_{\begin{subarray}{c}p^{\prime}\in\mathrm{L}(p+y)\\ p^{\prime}\in\mathrm{R}(p)\end{subarray}}\hat{\bm{Z}}_{p^{\prime}}^{\psi}\Big{]}\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p+y}^{\psi},$ (S11d)
where ${\rm R}(v)$ and $\mathrm{L}(v)$ denote all vertices to the right and to the left of $v$, respectively, and similarly for plaquettes. The mapping of Eq. (S11a) for horizontal processes is sketched in Fig. S2(c), while the mapping of Eq. (S11d) for vertical processes is sketched in Fig. S3(a)-(b).
@@
as reported in the main text, where we also used the hardcore-boson constraint to express the boson parities as $(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}=1-2\hat{b}_{v}^{\dagger}\hat{b}_{v}$ and we neglected constant energy shifts. The two-spin terms of Eq. (I) become
@@
The different stages of the mapping leading to Eq. (S17b) are sketched in S3 . For two-spin terms inducing fermion motion across the boundaries horizontally and vertically, respectively, the mapping yields
@@
Figure S3: Sketch of the different stages of the quasiparticle mapping for vertical fermion processes in the bulk [Eqs. (S11d) and (S17b)]. A two-spin term $\hat{Z}_{v,B}\hat{X}_{v,R}$ inducing tunnelling and pairing of the toric-code fermionic quasiparticles on plaquettes $p$ and $p+y$ is first mapped to a product of pseudospin operators $\hat{\bm{Z}}_{p}^{\psi}$, $\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p^{\prime}}^{\psi}$ and $\hat{\bm{Z}}_{v}^{e}$ (b). Then, such products are mapped to products of hardcore-boson operators on vertices ($\hat{b}_{v}$ and $\hat{b}_{v}^{\dagger}$, describing $e$ bosons) and Majorana-fermion operators on plaquettes ($\hat{\gamma}_{p}$ and $\hat{\gamma}_{p}^{\prime}$, related to fermionic operators $\hat{f}_{p}$ and $\hat{f}_{p}^{\dagger}$ for $\psi$ fermions), via a Jordan-Wigner mapping (c).
@@ II.1 Effective Hamiltonian
in terms of a time-independent Floquet Hamiltonian $\hat{H}_{F}^{[t_{0}]}$ and a time-periodic micromotion operator $\hat{U}_{F}(t,t_{0})=\hat{U}_{F}(t_{0}+T,t_{0})=e^{\hat{K}(t,t_{0})}$, generated by the ‘kick’ operator $\hat{K}(t,t_{0})$. The stroboscopic dynamics in steps of $T$ is then fully ruled by $\hat{H}_{F}^{[t_{0}]}$ according to
@@
The superscript in $\hat{H}_{F}^{[t_{0}]}$ makes explicit the parametric dependence on the initial time $t_{0}$ of the evolution and thus on the global phase of the drives at that time. As detailed in the following, for the driving scheme analyzed here, the stroboscopic generators at different $t_{0}$ (as well as the non-stroboscopic Floquet Hamiltonian [34]) match to leading order in high frequency and exhibit the desired topological properties. For this reason, in the main text we dropped the superscript on the Floquet Hamiltonian and denoted it with $\hat{H}_{F}$ (the impact of the phase of the drives on the non-Abelian vortex exchange phase is quantitatively analyzed below in the next section). Both $\hat{H}_{F}^{[t_{0}]}$ and $\hat{K}(t,t_{0})$ can be determined order by order in a high-frequency expansion in powers of $\omega^{-1}$ [34, 29]. The leading order is given by
@@
The leading order of the high-frequency expansion for the Floquet Hamiltonian $\hat{H}_{F}^{[t_{0}]\prime}$ is obtained by time averaging [Eq. (S31a)], which yields the time-independent part of Eq. (S35), namely
$\hat{H}_{F}^{[t_{0}]\prime}=\sum_{p}\Big{\{}-\mu\hat{f}_{p}^{\dagger}\hat{f}_{p}-\!\!\sum_{r\in\{x,y\}}\hat{\mathcal{S}}_{p}^{(r)}\big{(}J\hat{f}_{p}^{\dagger}\hat{f}_{p+r}\\ +\Delta e^{i\phi_{r}}\hat{f}_{p}\hat{f}_{p+r}+\mathrm{H.c.}\big{)}\Big{\}},$ (S36)
where we have defined the effective chemical potential $\mu=\omega/2-\mu_{\psi}$. In the original reference frame [before transforming through $\hat{R}(T)$], the Floquet Hamiltonian $\hat{H}_{F}^{[t_{0}]}$ is determined from the relation
$e^{-i\hat{H}_{F}^{[t_{0}]}T}$ $=\hat{U}(t_{0}+T,t_{0})=\hat{R}(T)\hat{V}(t_{0}+T,t_{0})$
$=e^{-i\pi\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}}e^{-i\hat{H}_{F}^{[t_{0}]\prime}T}$ (S37)
and the fact that $[\hat{R}(T),\hat{H}_{F}^{[t_{0}]\prime}]=0$ due to the conservation of the total fermion parity. The effect of $\hat{R}(T)$ is thus simply to shift the zero Bogoliubov–de Gennes quasienergy by $\pi/T$, that is half a Floquet-Brillouin zone. By choosing the gauge $\hat{f}_{p}\to e^{-i\phi_{x}/2}\hat{f}_{p}$, it becomes clear that the effective Hamiltonian of Eq. (S36) only depends, to leading order in high-frequency, on the phase-delay $\phi_{y}-\phi_{x}$ between the horizontal and vertical drives, and not on the global phase of the drives (the impact of the latter is analyzed in the following section). The kick operator $\hat{K}(t,t_{0})$, to leading order, is determined by Eq. (S31b), given $\hat{R}(t)$ and the Fourier components of $\hat{H}^{\prime}_{\psi}(t)$. The latter, from Eq. (S35), read

### v3

@@ Abstract
Floquet-Majorana modes.— We verify that this Floquet modulation indeed induces topological properties and yields Floquet-Majorana modes. We study these properties by numerically computing the Floquet Bogoliubov-de Gennes (BdG) Hamiltonian $\hat{H}_{F,\mathrm{BdG}}$, defined by $\hat{H}_{F}=(1/2)(\hat{\bm{f}}^{\dagger},\hat{\bm{f}})\hat{H}_{F,\mathrm{BdG}}(\hat{\bm{f}},\hat{\bm{f}}^{\dagger})^{T}$, as well as its quasienergy spectrum and ground state parity, from the driven model of Eq. (3[32]. The topological phase is expected for $|\mu|<4J$ and is characterized by a non-zero Chern number of the negative-energy fermion band [33, 34, 17]. Tuning the frequency $\omega$ to this topological region ($\mu\simeq-2J$), in the absence of $e$ vortices [$\mathcal{S}_{p}^{(r)}=1$], we find four nearly-degenerate high-frequency ‘ground’ states on a torus, one in each superselection sector corresponding to periodic or anti-periodic boundary conditions in $x$ and $y$ direction [33]. The topological degeneracy is approached exponentially with increasing system size, as shown in Fig. 2(a), and only three ground states have even fermion parity and are thus physical, signalling three Ising-anyon types $e,\mathds{1}_{\rm is},\psi$. On a cylinder, the quasienergy spectrum exhibits two edge modes [Fig. 2(b)], indicating the development of nontrivial topology and a non-zero Chern number.
@@
Introducing a pair of separated $e$ particles on the torus, approximate Majorana modes at near-zero quasienergy appear [Fig. 2(c)], which persist until $\omega$ is detuned out of the topological phase, as predicted by the time-averaged description of Eq. (5). For weak driving the gap remains open as the amplitude $J=\Delta$ increases, indicating that higher-order terms neglected in (5) do not fundamentally alter the nature of the phase in this regime. Their impact can be suppressed exponentially with decreasing $J/\omega$ [43, 44]. While generic periodically-driven systems are expected to reach “infinite-temperature” at long times due to energy absorption from the drive [45, 46], our model may evade such heating given that it maps to an integrable fermion system (and static bosons) [47, 48]. We verify in [32] that this is indeed the case at high frequency, also for large $e$-particle densities. Increasing the driving amplitudes beyond the high-frequency regime, while maintaining $\mu=-2J$, the fermions enter an anomalous Floquet phase [49, 50, 37]. Here the time-averaged description of Eq. (5) is no longer valid, but the system still exhibits Majorana modes, at both quasienergy zero and ±$\pm\pi/T$, see Fig. 2(c). This effect, only possible in driven systems [51, 52, 49, 53, 54, 55, 56, 57, 58], has been proposed as a means to realize non-Abelian braiding and quantum computation in 1D Kitaev quantum wires [59, 60, 61, 62].
@@
Non-Abelian exchange phases.— We analyze whether the exchange statistics of two $e$ particles (vortices) in the fermion topological phase is non-Abelian, as predicted in the presence of Majorana zero modes [4, 35]. Two vortices are exchanged explicitly in different fusion sectors following Levin-Wen’s protocol [63, 4, 64], sketched in Fig. 2(d), and Ref. [31]. The protocol computes the difference in the Berry phase accumulated by moving vortices along two paths $P$ and $P^{\prime}$, sharing the same set of positions but involving a different order of vortex moves. This guarantees that the resulting phase is only determined by the exchange statistics [63]. For each intermediate vortex configuration |$\ket{\vec{n}_{i}^{e}}$ along the paths, we numerically determine the corresponding even-parity ground state |$\ket{\Phi_{i}}$ of the Floquet-BdG Hamiltonian [32]. The Berry phase accumulated along the path $P$ is computed from the sequence of ground states as $\theta_{P}=\mathrm{Arg}\prod_{(i,i+1)\in P}\bra{\Phi_{i+1}}\hat{Z}^{f}_{i+1,i}\ket{\Phi_{i}}$. Here, $\hat{Z}^{f}_{i+1,i}$ is the fermionic part of the quasiparticle representation of the spin operator $\hat{Z}_{i}$ which converts the vortex configuration |$\ket{\vec{n}^{e}_{i}}$ into |$\ket{\vec{n}^{e}_{i+1}}$ by displacing one $e$ particle. Details about the evaluation of the matrix elements in $\theta_{P}$ are given in the SM [32]. The prediction for Ising anyons is that the exchange phase $R_{\mathds{1}_{\rm is}}^{\sigma\sigma}$ in the $\mathds{1}_{\rm is}$ fusion sector (where the vortices fuse to the vacuum) and $R_{\psi}^{\sigma\sigma}$ in the $\psi$ sector (where they fuse to a fermion $\psi$) differ by $\pi C/2$ for Chern number $C$, namely $R_{\mathds{1}_{\rm is}}^{\sigma\sigma}=e^{-i\pi{C}/8}$ and $R_{\psi}^{\sigma\sigma}=e^{3i\pi C/8}$ [4]— a signature of their non-Abelian nature [65]. The two sectors $\mathds{1}_{\rm is}$ and $\psi$ are selected by creating the $e$ pair on top of two different ground states, corresponding to doubly-antiperiodic and doubly-periodic fermion boundary conditions, respectively [31] [^1] In the antiperiodic case, the even-parity ground state is the quasiparticle vacuum, such that the pair of vortices is created from and fuses to the vacuum. For the periodic case, the BdG ground state has odd parity and thus the physical even-parity ground state is the first-excited state containing one BdG fermion. The vortex pair thus fuses to this fermion.. The counterclockwise exchange phases obtained are shown in Fig 2(e), for varying driving amplitude in the high-frequency regime and for different system size. The results converge rapidly with the size towards the prediction for Ising TO with $C=1$. We have thus shown that the $e$ vortices carrying Floquet-Majorana modes, arising in the driven toric code in the high-frequency regime, behave like non-Abelian Ising anyons. The stability of the exchange phases for increasing driving amplitude in this regime confirms that higher-order corrections to the time averaged Hamiltonian do not disrupt the topological phase, as anticipated. While the direct $e$ exchange probed here is a smoking-gun signature at zero temperature, fermion fractionalization can be detected at finite temperature, e.g., in the temperature dependence of Raman scattering intensities [67].
Conclusion.— We have shown that time-periodic driving of an Abelian-anyon system can induce non-Abelian topological order, using Kitaev’s toric code as the paradigmatic example of a large class of Abelian topological phases. Our findings suggest a potential path towards non-Abelian anyons in synthetic quantum systems, where Abelian phases have been realized [7, 8]. The model studied extends the range of potential candidates exhibiting 2D Floquet-Majorana physics beyond systems with intrinsic superconductivity or superfluidity [68, 69] and offers a flexible playground where key parameters such as quasiparticle motion and pairing processes can be independently controlled via the drive. The toric code, known to be closely related to quantum dimer models [70, 71, 72], has been recently shown to describe dimer liquids of Rydberg excitations constrained by Rydberg blockade [73, 74, 75], where Abelian spin-liquid behaviour has been observed [7]. This represents a promising setup to explore the ideas presented here, alongside adaptations of Floquet-engineering protocols for the toric code proposed in superconducting circuits [23, 25].
@@
Acknowledgments.— The author is grateful to André Eckardt, Nathan Harshman, Alexander Schnell, Isaac Tesfaye and Sandro Wimberger for helpful and inspiring discussions. The author also acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Research Unit FOR 2414 - project number No. 277974659.
@@
Non-Abelian Anyons in Periodically Driven Abelian Spin Liquids
@@ I Quasiparticle mapping
where $\Gamma_{1,v}$ is a path connecting vertices 1 and $v$, while $\tilde{\Gamma}_{1,p}$ connects the $1$st and the $p$th plaquette, with the specific path convention depicted in Fig. S1(a): $\Gamma_{1,v}$ descends vertically from the top-left vertex until the vertex row to which $v$ belongs and then continues horizontally until $v$; $\tilde{\Gamma}_{1,p}$ connects the top-left plaquette to the right-most one in the same row, descends vertically until the plaquette row to which $p$ belongs, and then horizontally to reach $p$. With this choice, it can be verified that spin anti-commutation rules are satisfied by the pseudospins, $\{\hat{\bm{X}}_{1}^{e}\hat{\bm{X}}_{v}^{e},\hat{\bm{Z}}_{v}^{e}\}=0$ and $\{\hat{\bm{X}}_{1}^{\psi}\hat{\bm{X}}_{p}^{\psi},\hat{\bm{Z}}_{p}^{\psi}\}=0$. The operators in Eq. (S10) are depicted in Fig. S2(a).
Using Eqs. (S6) and (S10), the two-spin driving terms of the form $\hat{X}_{i}\hat{Z}_{j}$ entering the drive Hamiltonian $\hat{H}_{d}(t)$ of Eq. (2) are mapped to pseudospin operators. The precise mapping depends on whether a term involves spins located on the boundaries, which are intersected by $\Gamma_{r}$ or $\tilde{\Gamma}_{r}$, or not [31]. We only report, as an example, the mapping for terms in the bulk, not intersecting the boundary, which reads
@@
×$\times\Big{[}\prod_{\begin{subarray}{c}p^{\prime}\in\mathrm{L}(p+y)\\ p^{\prime}\in\mathrm{R}(p)\end{subarray}}\hat{\bm{Z}}_{p^{\prime}}^{\psi}\Big{]}\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p+y}^{\psi},$ (S11b)
where ${\rm R}(v)$ and $\mathrm{L}(v)$ denote all vertices to the right and to the left of $v$, respectively, and similarly for plaquettes. The mapping of Eq. (S11a) for horizontal processes is sketched in Fig. S2(c), while the mapping of Eq. (S11b) for vertical processes is sketched in Fig. S3(a)-(b).
@@
as reported in the main text, where we also used the hardcore-boson constraint to express the boson parities as $(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}=1-2\hat{b}_{v}^{\dagger}\hat{b}_{v}$ and we neglected constant energy shifts. The two-spin terms of Eq. (S11) become
@@
The different stages of the mapping leading to Eq. (S17b) are sketched in Fig. S3. For two-spin terms inducing fermion motion across the boundaries horizontally and vertically, respectively, the mapping yields
@@
Figure S3: Sketch of the different stages of the quasiparticle mapping for vertical fermion processes in the bulk [Eqs. (S11b) and (S17b)]. A two-spin term $\hat{Z}_{v,B}\hat{X}_{v,R}$ inducing tunnelling and pairing of the toric-code fermionic quasiparticles on plaquettes $p$ and $p+y$ is first mapped to a product of pseudospin operators $\hat{\bm{Z}}_{p}^{\psi}$, $\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p^{\prime}}^{\psi}$ and $\hat{\bm{Z}}_{v}^{e}$ (b). Then, such products are mapped to products of hardcore-boson operators on vertices ($\hat{b}_{v}$ and $\hat{b}_{v}^{\dagger}$, describing $e$ bosons) and Majorana-fermion operators on plaquettes ($\hat{\gamma}_{p}$ and $\hat{\gamma}_{p}^{\prime}$, related to fermionic operators $\hat{f}_{p}$ and $\hat{f}_{p}^{\dagger}$ for $\psi$ fermions), via a Jordan-Wigner mapping (c).
@@ II.1 Effective Hamiltonian
in terms of a time-independent Floquet Hamiltonian $\hat{H}_{F}^{[t_{0}]}$ and a time-periodic micromotion operator $\hat{U}_{F}(t,t_{0})=\hat{U}_{F}(t+T,t_{0})=e^{\hat{K}(t,t_{0})}$, generated by the ‘kick’ operator $\hat{K}(t,t_{0})$. The stroboscopic dynamics in steps of $T$ is then fully ruled by $\hat{H}_{F}^{[t_{0}]}$ according to
@@
The superscript in $\hat{H}_{F}^{[t_{0}]}$ makes explicit the parametric dependence on the initial time $t_{0}$ of the evolution and thus on the global phase of the drives at that time. As detailed in the following, for the driving scheme analyzed here, the stroboscopic generators at different $t_{0}$ (as well as the non-stroboscopic Floquet Hamiltonian [41]) match to leading order in high frequency and exhibit the desired topological properties. For this reason, in the main text we dropped the superscript on the Floquet Hamiltonian and denoted it with $\hat{H}_{F}$ (the impact of the phase of the drives on the non-Abelian vortex exchange phase is quantitatively analyzed below in Sec. II.2). Both $\hat{H}_{F}^{[t_{0}]}$ and $\hat{K}(t,t_{0})$ can be determined order by order in a high-frequency expansion in powers of $\omega^{-1}$ [41, 36]. The leading order is given by
@@
The leading order of the high-frequency expansion for the Floquet Hamiltonian, $\hat{H}_{F}^{[t_{0}]\prime}$ is obtained by time averaging [Eq. (S31a)], which yields the time-independent part of Eq. (S35), namely
$\hat{H}_{\rm avg}^{[t_{0}]\prime}=\sum_{p}\Big{\{}-\mu\hat{f}_{p}^{\dagger}\hat{f}_{p}-\!\!\sum_{r\in\{x,y\}}\hat{\mathcal{S}}_{p}^{(r)}\big{(}J\hat{f}_{p}^{\dagger}\hat{f}_{p+r}\\ +\Delta e^{i\phi_{r}}\hat{f}_{p}\hat{f}_{p+r}+\mathrm{H.c.}\big{)}\Big{\}},$ (S36)
where we have defined the effective chemical potential $\mu=\omega/2-\mu_{\psi}$. This corresponds to the Hamiltonian of Eq. (4) of the main text. In the original reference frame [before transforming through $\hat{R}(T)$], the approximated Floquet Hamiltonian $\hat{H}_{\rm avg}^{[t_{0}]}$ is determined from the relation
$e^{-i\hat{H}_{\rm avg}^{[t_{0}]}T}$ $=\hat{U}(t_{0}+T,t_{0})=\hat{R}(T)\hat{V}(t_{0}+T,t_{0})$
$=e^{-i\pi\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}}e^{-i\hat{H}_{\rm avg}^{[t_{0}]\prime}T}$ (S37)
and the fact that $[\hat{R}(T),\hat{H}_{\rm avg}^{[t_{0}]\prime}]=0$ due to the conservation of the total fermion parity. The effect of $\hat{R}(T)$ is thus simply to shift the zero Bogoliubov–de Gennes quasienergy by $\pi/T$, that is half a Floquet-Brillouin zone. By choosing the gauge $\hat{f}_{p}\to e^{-i\phi_{x}/2}\hat{f}_{p}$, it becomes clear that the effective Hamiltonian of Eq. (S36) only depends, to leading order in high-frequency, on the phase-delay $\phi_{y}-\phi_{x}$ between the horizontal and vertical drives, and not on the global phase of the drives (the impact of the latter is analyzed in the following Sec. II.2). The kick operator $\hat{K}(t,t_{0})$, to leading order, is determined by Eq. (S31b), given $\hat{R}(t)$ and the Fourier components of $\hat{H}^{\prime}_{\psi}(t)$. The latter, from Eq. (S35), read